OpenStudy (jotopia34):

Someone please help me! I am trying to find a phase angle on my TI84. I know that Fi is equal to 4.25 from answer book. tan (Fi)= c1/c2 c1= -1 and c2= -1/2 soo tan(Fi) =2 to find Fi, I thought you do, arctan (2) = Fi, where Fi should be 4.25. But this doesnt happen. If I do tan(4.25), I get 2. But if I do arctan(2), I do not get 4.25. What am I doing wrong? Is there a setting I have wrong? (BTW, I am in radian mode on my TI84) thanx

OpenStudy (jotopia34):

also, if I input cos(2pi) I get 1, but if I do arccos(1), I get 0. Its like its not doing the inverse.

OpenStudy (anonymous):

fi = phi \[\text{phi}= \varphi\] fi is what giants say, when they smell the blood of an Englishman. For the cosine problem, \[ \cos(2 \pi) = \cos(0) \longrightarrow \cos(2 \pi n) = 1 \text{ for all integers}\] When you do arctan(2) you get ~ 1.107, right? 1.107+pi ~ 4.25 \[\tan(\theta + n\pi) = \tan \theta\]

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

What's the problem that you're trying to solve? That might give a little more insight into how you can figure out the way the book got 4.25. ^^

OpenStudy (jotopia34):

YAYYY, yes this is exactly what what wrong. I needed to add pi to it. You saved me hours of struggling. Thank you

OpenStudy (jotopia34):

oh, and Fo and Fum aren't so bad either ;)

OpenStudy (jotopia34):

I see now that you need to add pi, because 4.25 would be in the wrong quadrant for the values of C1 and C2 that are both negatives. Adding pi puts me into the right spot. Is that right?

OpenStudy (anonymous):

yeah, pho is awesome! ^_^ Also, yes; with a negative c1 and c2 it'd be in quadrant 3; so adding pi puts you up in quadrant 1 if that's where you needed to be. Granted, I don't know what your c1 and c2 stand for, but theoretically that's true! ^_^

OpenStudy (jotopia34):

Ph, not F, I now see this. ;) my C1 was -1 and C2 was -1/2

OpenStudy (jotopia34):

thks again.

OpenStudy (anonymous):

I was also talkin' bout some soup pho, not just driving the ph- train into the ground. fum doesn't get to go to that party 0_0 Welcome again, and holler if you have any more questions ^_^